# WindVE: Collaborative CPU-NPU Vector Embedding

### v2

@@
Abstract.
Retrieval-Augmented Generation (RAG) is a technology that enhances large language models (LLMs) by integrating information retrieval. It uses vector embedding as a key technology to map queries to relevant information, enriching the context for text generation. This integration empowers LLMs to generate more coherent and well-informed content. In the industry, inference services based on LLMs are highly sensitive to cost-performance ratio, prompting the need for improving hardware resource utilization in the inference service. Specifically, vector embedding and retrieval processes take up to 20% of the total latency. Therefore, optimizing the utilization of computational resources in vector embeddings is crucial for enhancing the cost-performance ratio of inference processes, which in turn boosts their product competitiveness.
In this paper, we analyze the deployment costs of vector embedding technology in inference services, propose a theoretical formula, and determine through the mathematical expression that increasing the capacity to process concurrent queries is the key to reducing the deployment costs of vector embeddings. Therefore, in this paper, we focus on improving the product’s capability to process concurrent queries. To optimize concurrency without sacrificing performance, we have designed a queue manager that adeptly offloads CPU peak queries. This manager utilizes a linear regression model to ascertain the optimal queue depths, a critical parameter that significantly influences the efficacy of the system. We further develop a system named WindVE that uses a CPU-NPU heterogeneous architecture to offload peak concurrent queries, which leverages the performance differences between the two processors to effectively manage traffic surges. Through experiments, we compare WindVE to the state-of-the-art vector embedding framework FlagEmbedding, and achieve a concurrency level up to 22.3% higher than the scheme without offloading. This means that the scheme of using CPU to offload peak concurrent queries will achieve up to 1.22$\times$ the throughput under the same hardware resource conditions; or, to meet the same performance requirements, CPU and NPU collaboration to offload peak concurrent queries can save up to 18.6% of the hardware deployment costs.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Symposium on Parallelism in Algorithms and Architectures 2025; July 28-August 1, 2025; Portland, Oregonisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Concurrent algorithms
1. Introduction
Figure 1. RAG workflow.
In today’s landscape, LLMs are increasingly gaining prominence and are extensively utilized across a multitude of industrial applications, including the Huawei O3 Knowledge Community (Huawei Technologies, 2024) and OpenAI’s ChatGPT (OpenAI, 2024). RAG is an innovative approach in LLM applications that enhances the accuracy (Reynolds and Corrigan, 2024) and prevents hallucination (Shuster et al., 2021) of LLMs by integrating information retrieval systems (Ke et al., 2024; Xu et al., 2023; Lála et al., 2023; Zhang et al., 2023; Thakur et al., 2024). As a core component of RAG, vector embedding uses numerical representations that convert text into high-dimensional vectors, capturing semantic meanings and relationships between words or sentences (Rau et al., 2024; Zhao et al., 2024b). In the context of RAG, vector embedding plays a vital role in both retrieval (Monir et al., 2024) and generation (Geng et al., 2022) phases of LLMs. Figure 1 depicts a typical RAG workflow. During retrieval phase, an embedding model is used to extract information related to the input query from an external database, ensuring that the generated text is informed and contextually appropriate. Relevant vector embeddings are then fused with the input query, guiding the generation process in a LLM to produce more refined and domain-specific outputs. In practical, the RAG routine is much more complicated and vector embedding is widely used in many modules of RAG such as query rewriting, intent identification, MRC, reranking and so on. In many LLM-based systems, the vector embedding service can be called tens of millions of times within a month by tens of millions of users, because the embedding process needs to be performed online at least once for each request, different from the offline computing in traditional recommendation system.
Nevertheless, from the business point of view, LLM inference services are highly sensitive to cost–performance ratio. The high cost associated with inference system deployment can lead to a high price for customers, ultimately resulting in a loss of product competitiveness. For example, the price of the lately released LLM inference service gpt-4o-mini is as low as \$0.150 / 1M input tokens (AI, 2024). Specifically, in modern industry applications, vector embedding and retrieval consume up to 20% of the total latency within a typical RAG-integrated LLM inference system (Gao et al., 2024). Therefore, companies must strive to optimize vector embeddings to ensure that they offer the best possible performance at a reasonable price.
@@ 1. Introduction
  • We design a queue manager to facilitate seamless CPU-NPU collaboration for vector embedding, coupled with a linear regression based estimator for ascertaining its critical parameters to maximize concurrency.
  • We develop WindVE, a flexible and scalable system framework for high-throughput offloading vector embedding with CPU-NPU collaboration.
  • We demonstrate WindVE incorporating CPU-NPU heterogeneous computing for peak query offloading in vector embedding applications. Our experimental results demonstrate that we have achieved 1.20$\times$ concurrency with SLO met.
The rest of this paper is organized as follows. Section 2 reviews the work focusing on improving the hardware efficient for vector embeddings. Section 3 introduces why system concurrency is crucial in reducing inference service costs by mathematical analysis, and derives the theoretical benefits of CPU peak query offloading. Section 4 presents the design overview, and the implementation details of our framework WindVE using peak queries offloading. In Section 5, we demonstrated how WindVE improves system concurrency and throughput. Finally, Section 6 concludes the paper.
2. Background
@@
2.1. Vector Embedding
A group of literature related to vector embedding aim to improve functionality: Ref (Li et al., 2024) proposed an embedding framework using a memory network to obtain cross-modal memory-enhanced embeddings; Ref (Chen et al., 2019) introduces a scalable algorithm for network embeddings to capture transitive relationships in a graph; Ref (Koopman et al., 2019) proposed an embedding method extending random projection to an average language vector to improve document embeddings discriminating ability; Ref (Chourasia et al., 2023) proposed an alignment-free embedding approach to generate a fixed-length feature vector representation from the raw sequencing reads without assembly; Ref (Lai et al., 2023) proposed adaptive embedding method to distinguish important features in dynamic data distribution in training; and Ref (Günther, 2018) exploited how to utilize word embeddings to augment and enrich queries in database management systems.
Another group of designs aim to improve the efficiency of vector embedding system: Ref (Hwang et al., 2012) employed nonlinear embeddings to project data points in computer vision area to low-dimensional space and proved to be more efficient than other embedding algorithms in finding the exact nearest neighbor based on the Euclidean distance; Ref (Weinberger et al., 2009; Yan et al., 2021; Shi et al., 2020) used hashing with collisions to compress embeddings to achieve higher processing efficiency; and Ref (Zhang et al., 2024) accelerated embeddings entailing only hash processes with one additional potential embedding lookup.
@@ 2.2. Inference Acceleration
A large group of works has contributed to LLM inference acceleration, such as ClossalAI (Li et al., 2023), vLLM (Kwon et al., 2023a), Triton (Triton, 2024), etc. ClossalAI provides an inference framework built upon Colossal-AI, to support parallelism for large-scale models, pre-built large models, engine encapsulation, and an online service system; vLLM gives an efficient and easy-to-use LLM inference and serving library with cutting-edge inference acceleration technologies such as paged attention (Kwon et al., 2023b), continuous batching, speculative decoding (Leviathan et al., 2023), etc; Triton is an inference server introduced by NVIDIA for standardized AI model deployment and execution empowered by dynamic batching, concurrent execution, optimal configuration, and streaming audio and video. These frameworks all implement inference acceleration using LLM characteristics, such as billions of parameters, parallelism among multiple cards, KVcache mechanism and prefill/decoding phases. In contrast, the embedding inference usually adopts a model smaller than 1B, without the need of KVcache and multi-card parallelism, as it can be well placed and processed by a single card or CPU cores. Also, the output of an embedding inference is a tensor instead of a word sequence, thus the decoding phase does not exist in this scenario. Therefore, existing LLM inference methods are not considered as major counterparts, as they cannot reproduce the same level of performance boost in embedding inference.
@@
For embedding model acceleration, FlagEmbedding (of Artificial Intelligence, 2024) is the state-of-the-art solution for vector embedding. Serving as an open-source library for enhancing sentiment representations of words or sentences, FlagEmbedding uses vector embedding to capture the sentiment relationships between tokens for more precise text understanding and information retrieval. However, it has some major drawbacks: it does not efficiently reduce high dimensions, causing sparse representations in high-dimensional space; it adds embedding in training phase, leading to high computational costs; it does not support processing in CPUs and NPUs at the same time. The first two shortcomings are introduced by FlagEmbedding’s embedding algorithm, whose improvement is not in scope of this paper, while overcoming the last fits the purpose of this paper, thus selected as the counterpart in our experiments in Section 5.
2.3. CPU-NPU Collaboration
@@
CPU-NPU heterogeneous computing has been widely used in applications such as power systems (Rodriguez et al., 2021), pollution dispersion (Fernandez et al., 2020), queue processing system (Rosenfeld et al., 2022), image processing (Naz et al., 2020), speech processing (Rakhimov and Ochilov, 2021), federated edge learning (Zeng et al., 2021), and pattern matching (Sanz et al., 2020). The detailed techniques of collaboration include coarse-grain thread-level parallelism across CPUs and NPUs/GPUs (Lee et al., 2012), CPU-NPU workload partitioning (Zhu et al., 2019; Zeng et al., 2021), and heterogeneous tensor or pipeline parallelism (Zhao et al., 2024a).
@@
These work all aim to reduce the responding latency of model inference, rather than the maximum concurrency, which is the focus of this paper. These two objectives are nearly contradictory, as increasing the batch size often leads to higher latency. Besides, the number of parameters in embedding models is usually lower than 1B (e.g., BGE 326M (Xiao et al., 2023), Jina 570M (Günther et al., 2023)). Thus, vector embedding can be run as an entire instance on CPU, eliminating the need for complex model slicing like Ref (Zhao et al., 2024a).
3. Motivation
@@
In this section, we first illustrate the motivation of this work through an analysis of why the system’s capacity to process the maximum number of concurrent queries is pivotal in diminishing the expenses associated with inference services. This analysis underpins our proposal to leverage CPUs that are already present in GPU/NPU server configurations to offload peak query loads, thereby enhancing concurrency. This strategy can significantly decrease the cost of request processing.
3.1. Why Concurrency Matters
@@
The latency of an inference service is defined as the total time taken from receiving a query to producing the corresponding output, the maximum concurrency of an inference service is the maximum amount of concurrent queries the system can process given the SLOs and memory constraints, and the throughput is the number of queries processed within a time window. In industry applications, the latency is a key performance indicator related to SLOs and user experience; the maximum concurrency indicates the system instantaneous capacity to process query bursts; and the throughput defines the average performance of query processing within the given time window, which can be affected by the actual incoming query distributions.
Figure 2. An illustration of query number changes in a day.
@@
(1) $t_{total}^{max}=t_{wait}^{max}+t_{proc}$
(2) $t_{total}=t_{wait}+t_{proc}$
@@
(3) $t_{wait}\leq t_{wait}^{max}=t_{total}^{max}-t_{proc}$
@@
(4) $n=\left\lfloor\frac{t_{total}^{max}-t_{proc}}{t_{proc}}\right\rfloor$
@@
(5) $\text{Cost}=\frac{\frac{N}{n}}{T}\cdot D\cdot P$
Where $D$ and $P$ are device number per instance and price per device, respectively. Service deployment according to the query number with a time window using Equation 5 is cost-saving because it utilizes hardware resources thoroughly. However, in typical industrial applications, online query numbers varies significantly during a day (as shown in Figure 2). Using an average processing ability $T$ to estimate the number of processors to deploy does not consider potential query bursts that excess the service capacity, which lead to overshoot in processing time. In the worst case, query bursts with number above average have risk of not meeting SLOs due to large waitting time.
@@
(6) $\text{Cost}=\frac{N_{peak}}{C}\cdot D\cdot P$
@@
  • maximum concurrency is widely used in serious business scenarios, such as fault maintenance and many popular Internet applications (e.g., Tiktok and rednote), which requires that the latency of each request is strictly not higher than the given border (called SLO).
@@
3.2. Opportunity: CPU Peak Query Offloading
According to the analysis in Section 3.1, NPUs/GPUs are prioritized to process queries to ensure performance quality. With the growth of query numbers, the maximum latency of processing in NPUs/GPUs increases, causing potential overtimes. Therefore, CPUs can be used to offload a small number of queries to mitigate timeout risks. The threshold to immigrate excessive queries to CPUs is the first hyperparameter to decide. The threshold $C_{NPU}^{max}$ must satisfy the following expressions:
(7) $t_{proc,NPU}^{C_{NPU}^{max}}\leq T$
(8) $t_{proc,NPU}^{C_{NPU}^{max}+1}>T$
@@ 3.2. Opportunity: CPU Peak Query Offloading
(9) $t_{proc,CPU}^{C_{CPU}^{max}}\leq T$
(10) $t_{proc,CPU}^{C_{CPU}^{max}+1}>T$
According to two deployment schemes mentioned in Equation 5 and Equation 6, we can analyze the theoretical saved costs of employing CPU-NPU collaboration method: if we deploy the system according to the maximum possible concurrent number, the system concurrency can be enlarged from $C_{NPU}$ to $C_{NPU}+C_{CPU}$, hence saving $\frac{C_{CPU}}{C_{CPU}+C_{NPU}}\times 100\%$ of deployment cost; if we use the average query numbers within a period of time, the system improves the average throughput up to $\frac{C_{CPU}}{C_{NPU}}\times 100\%$. Therefore, the maximum possible deployment cost decrement is $\frac{C_{CPU}}{C_{NPU}}\times 100\%$.
@@
(11) $t_{proc,CPU}^{1}>T\$
4. Peak Concurrent Queries Offloading
Figure 3. WindVE System Compared to existing work. (A) shows a typical GPU/NPU-based vector embedding system workflow, and (B) shows the workflow of WindVE.
In Section 3, the importance of system maximum concurrency has been illustrated and the theoretical advantages of offloading peak queries to the CPU has been elucidated. Drawing upon these insights, the design methodology of WindVE is expounded upon in this section.
4.1. System Design

### v4

@@
Abstract
Retrieval-Augmented Generation (RAG) enhances large language models (LLMs) using information retrieval and vector embeddings. This improves context for text generation but can impact cost-performance, as vector operations consume significant latency (up to 20%). Optimizing computational resources for vector embeddings is thus crucial for cost-effective inference.
This paper analyzes vector embedding deployment costs and finds that boosting concurrent query processing is key to reducing them. To achieve this without sacrificing performance, we designed a queue manager to offload CPU peak queries using linear regression for optimal queue depths. We also developed WindVE, a CPU-NPU system leveraging processor differences to manage traffic surges. Experiments show WindVE achieves 22.3% higher concurrency than a non-offloading scheme, translating to 1.22x throughput under the same hardware or up to 18.6% hardware cost savings for the same performance.
Index Terms:
@@ Abstract
I Introduction
Figure 1: RAG workflow.
In today’s landscape, LLMs are increasingly gaining prominence and are extensively utilized across a multitude of industrial applications, including the Huawei O3 Knowledge Community [1] and OpenAI’s ChatGPT [2]. RAG is an innovative approach in LLM applications that enhances the accuracy [3] and prevents hallucination [4] of LLMs by integrating information retrieval systems [5, 6, 7, 8, 9]. As a core component of RAG, vector embedding uses numerical representations that convert text into high-dimensional vectors, capturing semantic meanings and relationships between words or sentences [10, 11]. In the context of RAG, vector embedding plays a vital role in both retrieval [12] and generation [13] phases of LLMs. Figure 1 depicts a typical RAG workflow. During retrieval, an embedding model extracts relevant information from an external database. These embeddings are then fused with the query to guide the LLM, improving the generation’s context and specificity. Embeddings are also used in other RAG modules like query rewriting. Due to this online, per-request embedding requirement, vector embedding services face massive usage (millions of calls per month), unlike traditional offline systems.
Nevertheless, from the business point of view, LLM inference services are highly sensitive to cost–performance ratio. The high cost associated with inference system deployment can lead to a high price for customers, ultimately resulting in a loss of product competitiveness. For example, the price of the lately released LLM inference service gpt-4o-mini is as low as \$0.150 / 1M input tokens [14]. Specifically, in modern industry applications, vector embedding and retrieval consume up to 20% of the total latency within a typical RAG-integrated LLM inference system [15]. Therefore, companies must strive to optimize vector embeddings to ensure that they offer the best possible performance at a reasonable price.
@@ I Introduction
  • We designed a queue manager to facilitate seamless CPU-NPU collaboration for vector embedding, coupled with a linear regression based estimator for ascertaining its critical parameters to maximize concurrency.
  • We developed WindVE, a flexible and scalable system framework for high-throughput offloading vector embedding with CPU-NPU collaboration.
  • We demonstrated WindVE incorporating CPU-NPU heterogeneous computing for peak query offloading in vector embedding applications. Our experimental results demonstrate that we have achieved 1.20$\times$ concurrency with service-level objectives (SLOs) met.
This paper is organized as follows: Section II reviews hardware efficiency for vector embeddings. Section III analyzes why system concurrency reduces costs and explores CPU offloading benefits. Section IV details our WindVE framework for peak query offloading. Section V demonstrates WindVE’s improvements in concurrency and throughput. Section VI concludes.
II Background
@@
II-A Vector Embedding
A group of literature related to vector embedding aim to improve functionality: Ref [16] proposed an embedding framework using a memory network to obtain cross-modal memory-enhanced embeddings; Ref [17] introduces a scalable algorithm for network embeddings to capture transitive relationships in a graph; Ref [18] proposed an embedding method extending random projection to an average language vector to improve document embeddings discriminating ability; Ref [19] proposed an alignment-free embedding approach to generate a fixed-length feature vector representation from the raw sequencing reads without assembly; Ref [20] proposed adaptive embedding method to distinguish important features in dynamic data distribution in training; and Ref [21] exploited how to utilize word embeddings to augment and enrich queries in database management systems.
Another group of designs aim to improve the efficiency of vector embedding system: Ref [22] employed nonlinear embeddings to project data points in computer vision area to low-dimensional space and proved to be more efficient than other embedding algorithms in finding the exact nearest neighbor based on the Euclidean distance; Ref [23, 24, 25] used hashing with collisions to compress embeddings to achieve higher processing efficiency; and Ref [26] accelerated embeddings entailing only hash processes with one additional potential embedding lookup.
@@ II-B Inference Acceleration
A large group of works has contributed to LLM inference acceleration, such as ClossalAI [27], vLLM [28], Triton [29], etc. ClossalAI provides an inference framework built upon Colossal-AI, to support parallelism for large-scale models, pre-built large models, engine encapsulation, and an online service system; vLLM gives an efficient and easy-to-use LLM inference and serving library with cutting-edge inference acceleration technologies such as paged attention [30], continuous batching, speculative decoding [31], etc; Triton is an inference server introduced by NVIDIA for standardized AI model deployment and execution empowered by dynamic batching, concurrent execution, optimal configuration, and streaming audio and video. These frameworks all implement inference acceleration using LLM characteristics, such as billions of parameters, parallelism among multiple cards, KVCache mechanism and prefill/decoding phases. In contrast, the embedding inference usually adopts a model smaller than 1B, without the need of KVCache and multi-card parallelism, as it can be well placed and processed by a single card or CPU cores. Also, the output of an embedding inference is a tensor instead of a word sequence, thus the decoding phase does not exist in this scenario. Therefore, existing LLM inference methods are not considered as major counterparts, as they cannot reproduce the same level of performance boost in embedding inference.
@@
For embedding model acceleration, FlagEmbedding [32] is the state-of-the-art solution for vector embedding. Serving as an open-source library for enhancing sentiment representations of words or sentences, FlagEmbedding uses vector embedding to capture the sentiment relationships between tokens for more precise text understanding and information retrieval. However, it has some major drawbacks: it does not efficiently reduce high dimensions, causing sparse representations in high-dimensional space; it adds embedding in training phase, leading to high computational costs; it does not support processing in CPUs and NPUs at the same time. The first two shortcomings are introduced by FlagEmbedding’s embedding algorithm, whose improvement is not in scope of this paper, while overcoming the last fits the purpose of this paper, thus selected as the counterpart in our experiments in Section V.
II-C CPU-NPU Collaboration
@@
CPU-NPU heterogeneous computing has been widely used in applications such as power systems [33], pollution dispersion [34], queue processing system [35], image processing [36], speech processing [37], federated edge learning [38], and pattern matching [39]. The detailed techniques of collaboration include coarse-grain thread-level parallelism across CPUs and NPUs/GPUs [40], CPU-NPU workload partitioning [41, 38], and heterogeneous tensor or pipeline parallelism [42].
@@
These work all aim to reduce the responding latency of model inference, rather than the maximum concurrency, which is the focus of this paper. These two objectives are nearly contradictory, as increasing the batch size often leads to higher latency. Besides, the number of parameters in embedding models is usually lower than 1B (e.g., BGE 326M [43], Jina 570M [44]). Thus, vector embedding can be run as an entire instance on CPU, eliminating the need for complex model slicing like Ref [42].
III Motivation
@@
In this section, we first illustrate the motivation for this work by analyzing why the system’s capacity to process the maximum number of concurrent queries is pivotal in reducing the expenses associated with inference services. This analysis underpins our proposal to leverage CPUs that are already present in GPU/NPU server configurations to offload peak query loads, thereby enhancing concurrency. This strategy can significantly decrease the cost of request processing.
III-A Why Concurrency Matters
@@
The latency of an inference service is defined as the total time taken from receiving a query to producing the corresponding output. The maximum concurrency of an inference service is the maximum number of concurrent queries the system can process given the SLOs and memory constraints. The throughput is the number of queries processed within a time window. In industry applications, latency is a key performance indicator related to SLOs and user experience; maximum concurrency indicates the system’s instantaneous capacity to process query bursts; and throughput defines the average performance of query processing within the given time window, which can be affected by the actual incoming query distributions.
Figure 2: An illustration of query number changes in a day.
@@
$t_{total}^{max}=t_{wait}^{max}+t_{proc}$ (1)
$t_{total}=t_{wait}+t_{proc}$ (2)
@@
$t_{wait}\leq t_{wait}^{max}=t_{total}^{max}-t_{proc}$ (3)
@@
$n=\left\lfloor\frac{t_{total}^{max}-t_{proc}}{t_{proc}}\right\rfloor$ (4)
@@
$\text{Cost}=\frac{\frac{N}{n}}{T}\cdot D\cdot P$ (5)
Where $D$ and $P$ are device number per instance and price per device, respectively. Service deployment according to the query number with a time window using Equation 5 is cost-saving because it utilizes hardware resources thoroughly. However, in typical industrial applications, online query numbers varies significantly during a day (as shown in Figure 2). Using an average processing ability $T$ to estimate the number of processors to deploy does not consider potential query bursts that excess the service capacity, which lead to overshoot in processing time. In the worst case, query bursts with number above average have risk of not meeting SLOs due to large waiting time.
@@
$\text{Cost}=\frac{N_{peak}}{C}\cdot D\cdot P$ (6)
@@
  • maximum concurrency is widely used in serious business scenarios, such as fault maintenance and many popular Internet applications (e.g., Tiktok and rednote), which requires that the latency of each request is strictly not higher than the given border, also knows as SLOs.
@@
III-B Opportunity: CPU Peak Query Offloading
According to the analysis in Section III-A, NPUs/GPUs are prioritized to process queries to ensure performance quality. With the growth of query numbers, the maximum latency of processing in NPUs/GPUs increases, causing potential overtimes. Therefore, CPUs can be used to offload a small number of queries to mitigate timeout risks. The threshold to immigrate excessive queries to CPUs is the first hyperparameter to decide. The threshold $C_{NPU}^{max}$ must satisfy the following expressions:
$t_{proc,NPU}^{C_{NPU}^{max}}\leq T$ (7)
$t_{proc,NPU}^{C_{NPU}^{max}+1}>T$ (8)
@@ III-B Opportunity: CPU Peak Query Offloading
$t_{proc,CPU}^{C_{CPU}^{max}}\leq T$ (9)
$t_{proc,CPU}^{C_{CPU}^{max}+1}>T$ (10)
According to two deployment schemes mentioned in Equation 5 and Equation 6, we can analyze the theoretical saved costs of employing CPU-NPU collaboration method: if we deploy the system according to the maximum possible concurrent number, where both CPUs and NPUs have the maximum number of queries processed concurrently without any query waiting others to finish, there is no $t_{wait}$ for both CPUs and NPUs. Therefore, the system concurrency can be enlarged from $C_{NPU}$ to $C_{NPU}+C_{CPU}$, hence saving $\frac{C_{CPU}}{C_{CPU}+C_{NPU}}\times 100\%$ of deployment cost; if we use the average query numbers within a period of time, the system improves the average throughput up to $\frac{C_{CPU}}{C_{NPU}}\times 100\%$. Therefore, the maximum possible deployment cost decrement is $\frac{C_{CPU}}{C_{NPU}}\times 100\%$.
@@
$t_{proc,CPU}^{1}>T\$ (11)
IV Peak Concurrent Queries Offloading
Figure 3: WindVE System Compared to existing work. (A) shows a typical GPU/NPU-based vector embedding system workflow, and (B) shows the workflow of WindVE.
In Section III, the importance of system maximum concurrency has been illustrated and the theoretical advantages of offloading peak queries to the CPU has been elucidated. Drawing upon these insights, the design methodology of WindVE is expounded upon in this section.
IV-A System Design